OpenStudy (jhannybean):

if \(z=xe^{xy}\) and \(x=3st\) and \(y=t^2-s\) find \(\large \frac{\partial z}{\partial s}\) at \((3,0)\)

OpenStudy (jhannybean):

I'm just unsure where to plug in the point \((3,0)\)

OpenStudy (jhannybean):

@hartnn @iambatman

OpenStudy (xapproachesinfinity):

this asking to find slope at the point s=3! what's the problem here?

OpenStudy (jhannybean):

\(\large \frac{\partial z}{\partial s}= \frac{\partial z}{\partial x} \frac{\partial x}{\partial s} +\frac{\partial z}{\partial y} \frac{\partial y}{\partial s}\) \(\large\frac{\partial z}{\partial s}=(e^{xy}+xye^{xy})(3t) + (x^2e^{xy})(-1)\) from here... do I plug in the point (3,0) into x and y? or what do I do after this?

OpenStudy (xapproachesinfinity):

did you find the partial derivative

OpenStudy (jhannybean):

Yep :)

OpenStudy (xapproachesinfinity):

well (3,0) is related to s so you can't plug it in for x and y the question is asking the derivative of z at that point when s=3 I don't see any s in your partial derivative

OpenStudy (amistre64):

x=3st and y=t^2−s use the point to determine s and t if needed 3 = 3st 0 = t^2-s

zepdrix (zepdrix):

Yah you want to think of z as a function of x and y. x=3, y=0. I can certainly see the confusion though :) Since z is a function of s and t, not directly though.

OpenStudy (jhannybean):

Ah that way... so 3 = 3st 0 = t^2 - s st = 1 0 = (1/s)^2 - s t = 1/s ? 0 = 1/(s^2) - s 0 = s^-2 - s something along this path? :\

zepdrix (zepdrix):

Hmmm that seems like a longer way to go about it...\[\Large\rm 3=3st\]\[\Large\rm 0=t^2-s\qquad\implies\qquad s=t^2\]\[\Large\rm 3=3(t^2)t\qquad \implies\qquad t^3=1\]Should work out nicer that way, yes? :o

OpenStudy (amistre64):

yeah, or s=t=1 seems to work

OpenStudy (jhannybean):

I was embarrassed to say I forgot how to factor out the s in my method <.< and yeah, the alternate method does work better.

OpenStudy (jhannybean):

Ohh, I just remembered... 1/s^2 - s = 0 (1-s^3)/s^2 = 0 1-s^3 = 0 s^3 = 1 s = 1 Thank you guys so much! Haha.

OpenStudy (amistre64):

yw

zepdrix (zepdrix):

Ah there's the factoring c: hehe

OpenStudy (jhannybean):

It's not even the calculus that kills me in these problems, it's mostly the factoring...

OpenStudy (xapproachesinfinity):

I totally miss understood this! thanks to this guys i recaptured what's happening!

OpenStudy (xapproachesinfinity):

these*